Spectra of Neumann Laplacians in thin graph like domains

Peter Kuchment Texas A&M University, College Station

We discuss convergence of spectra of Neumann Laplace operators on thin domains to spectra of appropriate differential operators on graphs. The results are obtained jointly with H. Zeng.